Jefferson Lab HALL A Logbook for March 08 - All EXCEPT Automatic Entries

  • Main INDEX

    Monday

  • 229246 03/31/08 22:50 V. Sulkosky Target Ladder Encoder Positions
  • 229245 03/31/08 22:30 V. Sulkosky Target Screens
  • 229244 03/31/08 22:10 V. Sulkosky Preshower and Shower HV
  • 229243 03/31/08 22:07 V. Sulkosky Scaler map File
  • 229241 03/31/08 21:21 V. Sulkosky Spectrometers Set
  • 229238 03/31/08 19:59 A. Camsonne BB HV off
  • 229237 03/31/08 19:54 jpchen Alarms are set
  • 229236 03/31/08 19:45 jpchen taregt motion porblem earlier
  • 229235 03/31/08 19:44 jpchen Cryotarget cooldown completed. Coolant issue.
  • 229225 03/31/08 18:40 A. Camsonne BigBite thresholds in start of run
  • 229222 03/31/08 18:29 Jack Left Q3
  • 229217 03/31/08 16:44 Dezern Checklist
  • 229216 03/31/08 16:43 jpchen Switched to cold return. A few items fixed.
  • 229215 03/31/08 15:41 jpchen target configuration updated
  • 229212 03/31/08 14:42 A. Camsonne Compton center crystal cable
  • 229200 03/31/08 14:19 jpchen Cryotarget cooldown started
  • 229194 03/31/08 13:56 A. Camsonne Experiment status
  • 229184 03/31/08 12:41 A. Camsonne BB HV added to start of run
  • 229181 03/31/08 12:33 V. Sulkosky L-HRS Moved to 35.5 degrees
  • 229179 03/31/08 12:18 A. Camsonne Compton central crystal connected to fan in place of CMU splitter box
  • 229170 03/31/08 10:54 A. Camsonne Compton calorimeter in place
  • 229169 03/31/08 09:50 Higinbotham BigBite Angle Set to 54 Degrees
  • 229168 03/31/08 08:57 A. Camsonne Target moved to BeO target
  • 229135 03/31/08 00:53 Kalyan scaler gate on L-HRS/BigBite scalers
  • Sunday

  • 229103 03/30/08 21:46 Moffit NIM bin replaced.
  • 229091 03/30/08 21:17 E.Chudakov Moller polarimeter status (technical)
  • 229089 03/30/08 21:13 E.Chudakov Pion rejector status
  • 229076 03/30/08 20:29 Moffit Bad NIM bin (used for delta E trigger).. replacing.
  • 229066 03/30/08 17:46 Kalyan TWOARM Fastbus configuration
  • 229059 03/30/08 16:57 brads MWDC remote threshold control instructions
  • 229056 03/30/08 16:44 Moffit/Chira BigBite HV GUI moved to adaqfs
  • 229054 03/30/08 16:12 brads Target camera itself seems dead
  • 229053 03/30/08 14:26 Annand BB Scintillator HV fine tune
  • 229035 03/30/08 10:06 Annand BB Scintillator HV fine tune
  • 229030 03/30/08 09:20 A. Camsonne Update on beam delivery
  • 229029 03/30/08 09:07 A. Camsonne Target camera not working
  • Saturday

  • 229026 03/29/08 22:55 Moffit EDTM Simulated Coincidences
  • 229019 03/29/08 20:23 bogdanw Left arm VDC
  • 229014 03/29/08 19:04 E.Chudakov Moller polarimeter checkout
  • 229013 03/29/08 18:49 Kalyan Trigger timings wrt to T1
  • 229012 03/29/08 18:26 V. Sulkosky L-HRS Gas Cerenkov
  • 229010 03/29/08 18:09 E.Chudakov Beam HV crate moved to portserver port 03
  • 228994 03/29/08 16:03 folts checklist
  • 228989 03/29/08 14:50 Annand BigBite scintillator HV settings saved mar29_1972.set.
  • 228988 03/29/08 14:42 Annand Correct BB scintillator HV cable error.
  • 228978 03/29/08 13:35 Annand BigBite scint HV cabling error
  • 228977 03/29/08 13:13 A. Camsonne Distance target flange to BB flange
  • 228969 03/29/08 12:35 V. Sulkosky E04-007 Podd Detector Databases
  • 228960 03/29/08 11:19 V. Sulkosky L-HRS Gas Cerenkov HVs
  • 228955 03/29/08 11:00 V. Sulkosky R-HRS Helicity Restored
  • 228946 03/29/08 00:45 Moffit added entries to DecData database..
  • Friday

  • 228945 03/28/08 23:49 Moffit Zero suppression for BB ADCs (vme)
  • 228941 03/28/08 21:38 Moffit MWDC HV trip (twice)
  • 228934 03/28/08 19:23 Jack VDC covers removed
  • 228932 03/28/08 17:26 V. Sulkosky L-HRS HV GUI
  • 228931 03/28/08 17:19 V. Sulkosky L-HRS S1 and S2 Signals
  • 228919 03/28/08 14:26 R. Michaels T1,T2,EDTM to TDC
  • 228913 03/28/08 13:40 R. Michaels scaler & trigger notes
  • Thursday

  • 228902 03/27/08 22:09 jinhuang bbsoftw@adaqlx:~~/software/onlana->~~/e04007/onlana
  • 228900 03/27/08 17:56 Moffit Sweep completed. Hot checkout to begin
  • 228899 03/27/08 17:13 R. Michaels online DT estimate updated
  • 228888 03/27/08 15:05 Moffit Insane Rate Test, part 4
  • 228881 03/27/08 14:29 Moffit Insane rate test, part 3
  • 228878 03/27/08 14:23 Moffit Insane Rate Test, part 2.
  • 228876 03/27/08 14:15 Moffit Insane rate test. Part 1.
  • 228869 03/27/08 12:10 Higinbotham BigBite Field Measurements
  • Wednesday

  • 228860 03/26/08 23:53 jinhuang Update BigBiteLib @ adaqlx
  • 228856 03/26/08 19:56 V. Sulkosky Spot++
  • 228853 03/26/08 19:07 V. Sulkosky LHRS BPM/Raster ADC Cable
  • 228826 03/26/08 13:25 Moffit Re: BB ADC gates
  • 228823 03/26/08 12:52 Moffit BigBite ADC gate aligned.
  • Tuesday

  • 228816 03/25/08 23:03 Moffit Re: EDTM settings. ANOTHER correction.
  • 228792 03/25/08 17:30 R. Michaels pedrunL
  • 228778 03/25/08 15:24 bogdanw HRS-Left Detector and DAQ status
  • 228774 03/25/08 14:42 atrig Trigger-Download
  • 228773 03/25/08 14:12 R. Michaels HRS discrim -> 30
  • 228772 03/25/08 14:11 atrig Trigger-Download
  • 228770 03/25/08 12:00 R. Michaels 2 HRS DAQs - ready
  • 228768 03/25/08 11:46 atrig Trigger-Download
  • 228767 03/25/08 11:42 atrig Trigger-Download
  • 228766 03/25/08 11:41 atrig Trigger-Download
  • 228765 03/25/08 11:30 atrig Trigger-Download
  • 228764 03/25/08 11:28 atrig Trigger-Download
  • Monday

  • 228753 03/24/08 21:42 R. Michaels timing window on L-HRS
  • 228751 03/24/08 21:31 atrig Trigger-Download
  • 228749 03/24/08 21:15 atrig Trigger-Download
  • 228745 03/24/08 16:31 V. Sulkosky Low Pressure Alarm in Counting House
  • 228743 03/24/08 13:42 Moffit re: edtm settings.. two different momenta
  • 228742 03/24/08 11:54 V. Sulkosky Shower HV Card in Slot # 8
  • 228741 03/24/08 11:32 Jack BNC replaced on Right S2 6R PMT
  • Sunday

  • 228725 03/23/08 17:09 Moffit Re: EDTM settings... correction.
  • 228720 03/23/08 15:43 Moffit EDTM settings
  • Friday

  • 228717 03/21/08 20:28 V. Sulkosky R-HRS Gas Cerenkov and T2
  • Thursday

  • 228697 03/20/08 23:01 V. Sulkosky R-HRS Gas Cerenkov 2
  • 228692 03/20/08 14:43 R. Michaels 103.7 kHz clock and other synch chks
  • 228691 03/20/08 14:27 V. Sulkosky R-HRS Gas Cerenkov
  • 228690 03/20/08 14:22 atrig Trigger-Download
  • 228689 03/20/08 14:20 atrig Trigger-Download
  • 228688 03/20/08 13:36 atrig Trigger-Download
  • 228687 03/20/08 13:28 atrig Trigger-Download
  • 228686 03/20/08 13:18 atrig Trigger-Download
  • 228685 03/20/08 13:18 atrig Trigger-Download
  • 228684 03/20/08 13:16 atrig Trigger-Download
  • 228683 03/20/08 13:13 R. Michaels HV crate on, T4 definition
  • Wednesday

  • 228682 03/19/08 20:06 V. Sulkosky R-HRS S2 Recabled
  • 228681 03/19/08 18:38 Jack VDC gas system restored yesterday
  • 228680 03/19/08 16:46 R. Michaels new T7 on L-HRS
  • 228679 03/19/08 16:41 R. Michaels L-HRS HV problem
  • 228678 03/19/08 14:07 fansler Right Q3 He valves
  • Tuesday

  • 228677 03/18/08 20:40 V. Sulkosky Shower HV
  • 228671 03/18/08 17:30 R. Michaels inserting prescale factors
  • 228665 03/18/08 16:14 R. Michaels LeftHrs config
  • 228663 03/18/08 15:55 atrig Trigger-Download
  • 228662 03/18/08 15:53 atrig Trigger-Download
  • 228657 03/18/08 13:26 V. Sulkosky Helicity Signal Table 2
  • 228656 03/18/08 13:19 V. Sulkosky Redundant Helicity Signals in L-HRS
  • Monday

  • 228642 03/17/08 15:32 V. Sulkosky A2 decabled for Gas Cerenkov
  • 228641 03/17/08 12:13 R. Michaels EDTM pulser
  • Sunday

  • 228640 03/16/08 23:00 Moffit Pi0 Work this weekend
  • 228639 03/16/08 14:30 V. Sulkosky Shower HV Channels 61 to 80
  • 228638 03/16/08 14:21 V. Sulkosky Shower HV / A2 HV
  • Saturday

  • 228637 03/15/08 12:23 R. Michaels TWOARM config
  • Friday

  • 228636 03/14/08 15:49 R. Michaels L-HRS DAQ work
  • 228597 03/14/08 09:37 A. Camsonne Pedestal subtraction off
  • Wednesday

  • 228596 03/12/08 18:45 A. Camsonne NaI ADC removed
  • 228595 03/12/08 10:32 Jack VDC gas mixng system is off
  • Tuesday

  • 228594 03/11/08 17:42 Moffit ODH Alarm
  • 228588 03/11/08 15:54 R. Michaels R-HRS DAQ
  • 228587 03/11/08 15:52 atrig Trigger-Download
  • 228585 03/11/08 15:36 atrig Trigger-Download
  • 228584 03/11/08 15:24 atrig Trigger-Download
  • 228579 03/11/08 11:53 R. Michaels duplicate files
  • Monday

  • 228578 03/10/08 17:52 Jack FPP Gas System
  • 228577 03/10/08 17:51 Jack Right Arm Detector Stack
  • Friday

  • 228576 03/07/08 13:39 fansler Left/Right Dipole He levels converted to 0-10V
  • Wednesday

  • 228575 03/05/08 11:34 R. Michaels DAQ work
  • Tuesday

  • 228571 03/04/08 09:23 Jack Detector work yesterday
  • Saturday

  • 228570 03/01/08 20:11 Moffit Pi0 Work Today